# 4D Scaffold Gaussian Splatting for Efficient Dynamic Scene Reconstruction

### v1

@@ Appendix G More results
4DGaussians 28.44 0.919 0.060 33.10 0.953 0.042 33.12 0.954 0.044
E-D3DGS 29.10 0.931 0.042 32.96 0.956 0.034 33.57 0.958 0.034
4DGS 28.63 0.918 0.071 33.54 0.956 0.039 34.18 0.959 0.038
STG 28.26 0.915 0.074 33.07 0.958 0.038 33.38 0.959 0.039
Ours 28.80 0.923 0.060 33.36 0.955 0.036 33.35 0.957 0.036
Model Metric flame_salmon flame_steak sear_steak
@@
4DGaussians 28.80 0.926 0.060 33.55 0.961 0.032 34.02 0.963 0.031
E-D3DGS 29.61 0.936 0.038 33.57 0.964 0.028 33.45 0.963 0.030
4DGS 29.25 0.929 0.057 33.90 0.961 0.037 33.37 0.960 0.040
STG 29.44 0.925 0.067 33.82 0.965 0.031 33.77 0.966 0.031
Ours 29.20 0.928 0.054 33.73 0.960 0.030 33.73 0.960 0.031
Table S4: Per-scene quantitative results on the Technicolor dataset.
Model Metric Birthday Fabien Painter
[…]
4DGaussians 28.03 0.862 0.156 33.36 0.865 0.185 34.52 0.899 0.138
E-D3DGS 33.34 0.951 0.037 34.45 0.875 0.147 36.63 0.923 0.097
4DGS 31.81 0.937 0.040 35.02 0.884 0.144 35.71 0.923 0.104
STG 31.88 0.947 0.028 35.79 0.900 0.115 35.53 0.925 0.095
Ours 32.86 0.952 0.024 35.32 0.889 0.140 37.35 0.941 0.069
Model Metric Theater Train
[…]
4DGaussians 28.67 0.835 0.195 23.54 0.756 0.206
E-D3DGS 30.64 0.864 0.147 31.84 0.922 0.074
4DGS 31.94 0.875 0.143 32.25 0.932 0.046
STG 31.25 0.880 0.133 32.58 0.938 0.039
Ours 31.65 0.891 0.106 32.69 0.941 0.032
Figure S6: Qualitative comparisons on the N3DV dataset.
Figure S7: Qualitative comparisons on the Technicolor dataset.

### v2

@@ Conclusion
coffee_martini 4DGaussian 23.83 0.821 0.182 - 28.44 0.919 0.060 193K 58 MB
E-D3DGS 26.05 0.901 0.096 - 29.10 0.931 0.042 263K 95 MB
Grid4D 25.76 0.874 0.139 - 28.69 0.919 0.058 153K 36 MB
STG 25.01 0.884 0.087 63K 27.45 0.912 0.070 481K 70 MB
4DGS 26.90 0.920 0.065 4229K 28.63 0.918 0.071 4293K 7978 MB
Ex4DGS 25.01 0.880 0.103 48K 28.79 0.918 0.067 340K 130 MB
Ours 27.07 0.921 0.053 447K 28.80 0.923 0.060 729K 116 MB
cook_spinach 4DGaussian 24.99 0.783 0.202 - 33.10 0.953 0.042 196K 58 MB
E-D3DGS 26.08 0.824 0.140 - 32.96 0.956 0.034 140K 51 MB
Grid4D 25.88 0.830 0.163 - 32.90 0.957 0.035 231K 55 MB
STG 25.02 0.797 0.195 40K 32.25 0.948 0.049 430K 62 MB
4DGS 27.32 0.879 0.088 3514K 33.54 0.956 0.039 3552K 6599 MB
Ex4DGS 26.10 0.844 0.137 60K 33.22 0.955 0.041 247K 118 MB
Ours 27.39 0.878 0.093 584K 33.36 0.955 0.036 782K 198 MB
cut_roasted_beef 4DGaussian 27.19 0.834 0.183 - 33.12 0.954 0.044 184K 55 MB
E-D3DGS 29.22 0.891 0.123 - 33.57 0.958 0.034 143K 52 MB
Grid4D 28.74 0.886 0.133 - 33.61 0.958 0.034 229K 54 MB
STG 27.87 0.864 0.150 38K 32.73 0.951 0.048 383K 56 MB
4DGS 31.21 0.936 0.060 3010K 34.18 0.959 0.038 3047K 5661 MB
Ex4DGS 29.56 0.904 0.099 64K 33.72 0.957 0.039 249K 123 MB
Ours 31.80 0.953 0.043 745K 33.35 0.957 0.036 916K 170 MB
flame_salmon 4DGaussian 21.98 0.832 0.163 - 28.80 0.926 0.060 197K 58 MB
E-D3DGS 24.29 0.890 0.104 - 29.61 0.936 0.038 264K 96 MB
Grid4D 24.52 0.888 0.113 - 29.86 0.930 0.051 158K 37 MB
STG 22.09 0.848 0.120 62K 28.27 0.916 0.065 562K 81 MB
4DGS 24.43 0.900 0.078 4774K 29.25 0.929 0.057 4782K 8886 MB
Ex4DGS 22.75 0.860 0.128 49K 28.78 0.924 0.064 330K 128 MB
Ours 24.82 0.911 0.059 473K 29.20 0.928 0.054 896K 162 MB
flame_steak 4DGaussian 25.43 0.855 0.134 - 33.55 0.961 0.032 186K 56 MB
E-D3DGS 26.54 0.891 0.101 - 33.57 0.964 0.028 135K 50 MB
Grid4D 25.77 0.880 0.117 - 32.98 0.963 0.029 214K 50 MB
STG 25.95 0.868 0.122 31K 33.35 0.958 0.037 373K 54 MB
4DGS 26.74 0.893 0.081 2439K 33.90 0.961 0.037 2451K 4553 MB
Ex4DGS 27.03 0.905 0.077 47K 33.90 0.962 0.032 226K 101 MB
Ours 27.84 0.922 0.049 500K 33.43 0.962 0.030 734K 142 MB
sear_steak 4DGaussian 28.59 0.876 0.127 - 34.02 0.963 0.031 198K 59 MB
E-D3DGS 29.36 0.906 0.107 - 33.45 0.963 0.030 135K 49 MB
Grid4D 29.27 0.905 0.109 - 33.49 0.965 0.028 228K 54 MB
STG 29.09 0.901 0.087 32K 33.40 0.960 0.034 371K 54 MB
4DGS 29.31 0.913 0.082 1866K 33.37 0.960 0.040 1880K 3492 MB
Ex4DGS 30.52 0.927 0.063 41K 33.68 0.960 0.034 221K 94 MB
Ours 31.44 0.945 0.045 448K 33.52 0.960 0.030 595K 109 MB
Table S4: Quantitative results on the N3DV dataset.
Dynamic region Full region
@@
Birthday 4DGaussian 25.14 0.838 0.150 - 28.03 0.862 0.156 297K 79 MB
E-D3DGS 29.10 0.948 0.041 - 33.34 0.951 0.037 231K 84 MB
STG 28.98 0.950 0.037 123K 32.20 0.947 0.029 302K 44 MB
C3DGS 28.03 0.933 0.054 126K 31.37 0.938 0.039 177K 26 MB
4DGS 29.27 0.953 0.032 7233K 31.72 0.936 0.042 7366K 13684 MB
Ex4DGS 29.07 0.949 0.033 169K 32.25 0.942 0.031 460K 162 MB
Ours 29.51 0.956 0.026 395K 32.86 0.952 0.024 623K 183 MB
Fabien 4DGaussian 28.54 0.807 0.283 - 33.36 0.865 0.185 224K 61 MB
E-D3DGS 30.23 0.859 0.214 - 34.45 0.875 0.147 173K 63 MB
STG 29.90 0.848 0.240 26K 34.69 0.876 0.166 72K 10 MB
C3DGS 29.54 0.836 0.260 30K 34.36 0.872 0.174 44K 6 MB
4DGS 31.44 0.892 0.170 2304K 35.02 0.884 0.144 2327K 4322 MB
Ex4DGS 31.33 0.890 0.155 116K 34.98 0.889 0.131 202K 83 MB
Ours 31.44 0.896 0.151 380K 35.15 0.895 0.124 398K 136 MB
Painter 4DGaussian 27.31 0.819 0.187 - 34.52 0.899 0.138 184K 51 MB
E-D3DGS 31.60 0.916 0.124 - 36.63 0.923 0.097 206K 75 MB
STG 32.10 0.925 0.120 61K 36.66 0.923 0.097 140K 20 MB
C3DGS 30.39 0.903 0.145 61K 35.55 0.913 0.113 75K 11 MB
4DGS 33.51 0.946 0.097 4945K 35.71 0.923 0.104 5015K 9316 MB
Ex4DGS 33.33 0.945 0.096 151K 36.62 0.932 0.091 255K 106 MB
Ours 35.19 0.960 0.075 1729K 37.05 0.939 0.073 1834K 465 MB
Train 4DGaussian 17.11 0.470 0.398 - 23.54 0.756 0.206 332K 88 MB
E-D3DGS 28.48 0.914 0.068 - 31.84 0.922 0.074 216K 78 MB
STG 26.56 0.896 0.110 101K 32.20 0.935 0.044 314K 46 MB
C3DGS 25.23 0.855 0.174 138K 30.79 0.910 0.071 193K 28 MB
4DGS 33.09 0.956 0.034 6468K 32.15 0.930 0.047 6980K 12967 MB
Ex4DGS 29.97 0.943 0.058 48K 31.40 0.927 0.055 816K 216 MB
Ours 32.42 0.964 0.026 2339K 32.86 0.946 0.032 2433K 275 MB
Theater 4DGaussian 21.86 0.612 0.381 - 28.67 0.835 0.195 302K 80 MB
E-D3DGS 27.56 0.794 0.291 - 30.64 0.864 0.147 237K 86 MB
STG 24.19 0.727 0.360 45K 30.92 0.878 0.143 188K 27 MB
C3DGS 22.85 0.690 0.411 81K 30.54 0.872 0.155 123K 18 MB
4DGS 32.25 0.904 0.132 7036K 31.90 0.877 0.136 7110K 13208 MB
Ex4DGS 29.21 0.862 0.204 119K 31.74 0.879 0.131 399K 132 MB
Ours 31.12 0.885 0.188 980K 32.65 0.893 0.106 1073K 332 MB
Table S5: Quantitative results on the Technicolor dataset.
Figure S7: Qualitative comparisons on the N3DV dataset with full region.
Figure S8: Qualitative comparisons on the Technicolor dataset with full region.